# Water-based Quantum Dots Liquid Scintillator for Particle Physics

### v1

@@ 6 Conclusions
In this article, we investigate the potential of water-based quantum dots liquid scintillator for particle physics application. We first performed a phase transfer to CdS/ZnS QDs from hydrophobic phase to hydrophillic phase. Absorbance and emission peaks for CdS QDs are $\sim 470$ nm and $\sim 460$ nm for both in toluene and in water. These values fit in the range provided by manufacturer, and we confirmed optical properties are mostly preserved in water-based quantum dots.
Photo luminescence quantum efficiency of the 2023 WbQD sample is measured to be $9.5\pm 0.5\%$. The value is lower than manufacture number, suggesting the loss due to phase transfer. Measured PLQY of other samples varies, and values go up or down over time and further study is needed to understand.
@@
In conclusion, water-based qunatum dots liquid scintillator is a promising new technology for particle physics. We have a special attention for a water solution of quantum dots to detect inverse beta decay processes from nuclear fission reactor neutrinos, where numerous programs are running all over the world [22]. Although liquid scintillator detectors [61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72] are popular base design, other technologies such as solid scintillator [73, 74, 75, 76] are also explored on top of water based detectors [16, 17, 18, 19, 20, 77, 78]. Furthermore, reactor neutrino detection from neutrino-electron elastic scattering [79, 80] and neutrino-nucleus elastic scattering are established [81, 82].

### v2

@@ 6 Conclusions
In this article, we investigate the potential of water-based quantum dots liquid scintillator for particle physics application. We first performed a phase transfer to CdS/ZnS QDs from hydrophobic phase to hydrophilic phase. Absorbance and emission peaks for CdS QDs are $\sim 470$ nm and $\sim 460$ nm for both in toluene and in water. These values fit in the range provided by the manufacturer, and we confirmed optical properties are mostly preserved in water-based quantum dots.
Photoluminescence quantum efficiency of the 2023 WbQD sample is measured to be $9.5\pm 0.5\%$. The value is lower than that provided by the manufacture, suggesting the loss due to phase transfer. Measured PLQY of other samples varies, and values go up or down over time and further study is needed to understand.
@@
In conclusion, water-based quantum dots liquid scintillator is a promising new technology for particle physics. We have a special attention for a water solution of quantum dots to detect inverse beta decay processes from nuclear fission reactor neutrinos, where numerous programs are running all over the world [24]. Although liquid scintillator detectors [63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74] are popular base designs, other technologies such as solid scintillator [75, 76, 77, 78] are also explored on top of water based detectors [16, 17, 18, 19, 20, 21, 22, 79, 80]. Furthermore, reactor neutrino detection from neutrino-electron elastic scattering [81, 82] and neutrino-nucleus elastic scattering are established [83, 84].